a cyclist traveled to her destination at an average rate of 15 mph. By traveling 3 mph faster, she took 30 minutes less to return. what distance did she travel each way?
Answer by stanbon(75887)

a cyclist traveled to her destination at an average rate of 15 mph. By traveling 3 mph faster, she took 30 minutes less to return. what distance did she travel each way?
----
Slow speed DATA:
rate = 15 mph; distance = x miles ; time = d/r = x/15 hrs
----
Faster speed DATA:
rate = 18 mph; distance = x miles ; time = x/18 hrs
----
Equation:
slow time - faster time = 1/2 hr
x/15 - x/18 = 1/2
-----
Multiply thru by 90 to get:
6x - 5x = 45
x = 45 miles (distance she traveled each way)
====================
Cheers,
Stan H.
==============